OpenStudy (anonymous):

What is the Least Common Multiple (LCM) of the two monomials: 450x2y5 and 3,000x4y3? Select one: 2xy 30xy 150x^2y^3 9,000x^4y^5

OpenStudy (anonymous):

find the LCM of 450 & 3000

OpenStudy (anonymous):

2xy?

OpenStudy (anonymous):

 LCM is 10 * 5 * 3 * 3 * 20

OpenStudy (anonymous):

9000

OpenStudy (anonymous):

yes For LCM...you take the highest value of the exponents

OpenStudy (anonymous):

oh wow can you help me with some other problems?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which of the following inequalities is true for all real values of x? (Hint: Try plugging in some positive and negative values for x to help see which one works all the time). Select one: x^3 ≥ x^2 3x^2 ≥ 2x^3 (2x)^2 ≥ 3x^2 3(x – 2)^2 ≥ 3x^2 – 2

OpenStudy (anonymous):

what have you tried

OpenStudy (anonymous):

none because I don't get it

OpenStudy (anonymous):

Test the situations: x^3> x^2 Try x = -2 -2^3 > -2^2 -8>4 no...so it's not a

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

3x^2 > 3x^3 let x = 2 3*2^2 > 3*2^3 3*4 > 3*8 nope...not today

OpenStudy (anonymous):

(2x)^2 > 3x^2 let x = 4 (2*4)^2 > 3*4^2 8^2 > 3*16 16> 48 Nope

OpenStudy (anonymous):

3(x – 2)^2 ≥ 3x^2 – 2 yes

OpenStudy (anonymous):

Test a number to be sure

OpenStudy (anonymous):

ok